A car drives at a speed of 26 m/s along a road parallel to a railroad track. A train traveling at 13 m/s sounds a horn that vibrates at 300 Hz.If the train and car are moving toward each other, what frequency of sound is heard by a person in the car?If the train and car are moving away from each other, what frequency of sound is heard in the car?
Answer #1

F1 = F*(V+Vo)/(V-Vs) when they will move towards each other

where Vo is speed of observer and Vs is speed of source .

=> F1 = 300*(343+26)/(343-12) = 334.44 Hz ( i am takin speed of sound to be 343 m/s)

b) F2 = 300*(343-26)/(343+12) = 267.88 Hz